# Gradient Existence and Energy Finiteness of Local Minimizers in the Wasserstein $L^{\infty}$ Topology for Binary-Star Systems

### v1

@@ 1 Introduction
A fundamental hydrodynamic model for Newtonian stellar dynamics is given by the Euler–Poisson system, describing the evolution of an isolated self‑gravitating fluid.
$\begin{split}\partial_{t}\rho+\nabla\cdot\left({\rho v}\right)&=0,\\ \rho\partial_{t}v+\rho\left({v\cdot\nabla}\right)v+\nabla P(\rho)&=\rho\nabla V,\\ \Delta V&=-4\pi\rho.\end{split}$ (EP)
where $\rho(x,t)\geq 0$ is the density of fluids at position $x\in\mathbb{R}^{3}$ and time $t\geq 0$, $v(x,t)\in\mathbb{R}^{3}$ is the velocity, and $V(x,t)\in\mathbb{R}$ is the gravitational potential. Pressure $P(\rho)$ depends on the density only, with some appropriate assumptions similar to McCann’s [31].
The existence of single-star and multi-star systems has been established by variational methods (see [4, 5, 3, 10, 11, 26, 31]) and perturbative approaches (see [21, 22, 19, 36, 37, 1, 27]).
As Jang and Seok discussed in [23], in the case of a uniformly rotating system, where we simply consider only the orbital revolution around each other and ignore the rotations of objects around their own axes, the Euler–Poisson equations reduce to a single equation:
$\omega^{2}\tilde{\rho}\left(x\right)P_{12}\left(x\right)+\nabla P\left(\tilde{\rho}\left(x\right)\right)-\tilde{\rho}\left(x\right)\left(\nabla V_{\tilde{\rho}}\left(x\right)\right)=0$ (EP’)
where $\omega>0$ is angular velocity, $P_{12}\left(x_{1},x_{2},x_{3}\right)=(x_{1},x_{2},0)$ , $\tilde{\rho}$ is a density function (in general with compact support), and $V_{\tilde{\rho}}(x)={\int{\frac{\tilde{\rho}(y)}{\left|{y-x}\right|}d^{3}y}}$. And the discussions in subsection 2.2 below shows that once $\tilde{\rho}$ satisfies the reduced Euler-Poisson equation (EP’), then $\rho\left({t,y}\right)=\tilde{\rho}\left(R_{-\omega t}y\right)$, $v\left({t,y}\right)=\omega\left({-y_{2},y_{1},0}\right)^{T}$ satisfies the original system (EP), where $R_{-\omega t}y$ is the rotation map given in Definition 2.1.
In [31], McCann used a variational method to construct uniformly rotating binary‑star solutions. His binary stars exhibit a dual characterization: as Hamiltonian and energy minimizers under conservative constraints, and as perturbations of simpler objects like non-rotating Lane-Emden stars or relative equilibria for point masses (e.g., [23, Section 3.2]). This characterization is essential for showing that the support of the energy constrained minimizer $\rho$ is contained within a ball of a certain radius, making $\rho$ a local minimizer with respect to the topology induced by Wasserstein $L^{\infty}$ (denoted by $W^{\infty}$ hereafter) distance. As discussed in Theorem 2.25 (see also [31, Theorem 2.1]), this ensures the existence of $\nabla P(\rho)$ in $\mathbb{R}^{3}$ and validates (EP’) in the whole space. However, McCann does not thoroughly address the existence of $\nabla P(\rho)$. While McCann provides valuable insights into the $W^{\infty}$ distance, further exploration is warranted—particularly why the local minimizer in this topology has finite energy, enabling the existence of the variational derivatives. These are what we aim to refine in this paper.
@@
    Gradient Existence and Equation Derivation: We establish the existence of $\nabla P(\rho)$ for local minimizer $\rho$, which enables a rigorous transition from the Euler-Lagrange equation to the Euler-Poisson equation. The result is formalized in Theorem 2.25 and proved in Section 3.
@@
    Energy Finiteness and Comparison of Topologies: We furthermore establish the finiteness of energy and the variational derivatives for local minimizers in Lemma 5.7. We also discuss results in the alternative topology inherited from the topological vector space, where we show finite‑energy local minimizers do not exist (Proposition 5.15), yet weak local minimizers with infinite energy can be found (Proposition 5.23).
Part of the results presented in this paper are based on the author’s Master’s thesis [12]. Beyond their intrinsic mathematical interest, these results serve as preliminary findings and will be used in the author’s companion paper [13] on star–planet systems, which also originates from the same thesis. In [13], they serve as an alternative to direct citations of McCann’s conclusions [31], providing a more precise and rigorous foundation for the arguments therein.
@@ 2 McCann’s Construction and Results of 2-Body Systems
In this section we review McCann’s construction and results of 2-Body systems [31]. In the first subsection, we introduce some basic settings, notations, assumptions and problem settings such as variational formulation. In the second subsection, we focus on explaining why the general energy form $E(\rho,v)$ can be replaced by the energy with uniform rotation $E_{J}(\rho)$. In the third subsection, we present some of McCann’s nice conclusions.
@@ 2.1 Notations and problem setting
    k{k} dimensional open ball: $B_{R}(a):=\left\{x\in\mathbb{R}^{k}\mid|x-a\mid<R\right\}$ for given $a\in\mathbb{R}^{k},R\in\mathbb{R}$.
@@
    The projection operator of $x$ to the $x_{1}x_{2}$ plane: $P_{12}(x)=P_{12}\left(x_{1},x_{2},x_{3}\right):=\left(x_{1},x_{2},0\right)$.
@@
    $r(x):=\left(\langle x,x\rangle_{2}\right)^{\frac{1}{2}}=x_{1}^{2}+x_{2}^{2}$.
@@
    Rotation map: R_θ:=(cosθ-sinθ0sinθcosθ0001)
  6. (vi)
    Nonnegative (Positive) part $\lambda]_{+}:=\max\{\lambda,0\}$.
  7. (vii)
@@
  8. (viii)
    Given $\delta>0$ the $\delta$-neighbourhood of $\Omega\subset\mathbb{R}^{3}$ to be the set Ω+B_δ(0):=⋃_y ∈Ω{x ∈R^3∣— x-y —¡δ}
  9. (ix)
    $\mu$: Lebesgue measure.
The state of a fluid may be represented by its mass density $\rho(x)\geq 0$ and velocity vector field $v(x)$ on $\mathbb{R}^{3}$. The fluid interacts with itself through Newtonian gravity hence we need to consider gravitational interaction energy, which will be given later. Moreover, to define internal energy, we follow Auchmuty and Beals’s assumptions [5] and first consider a general form of the pressure $P(\rho)$ as the following:
@@
    $P:[0,\infty)\rightarrow[0,\infty)$ continuous and strictly increasing;
@@
    $\lim\limits_{\rho\rightarrow 0}P(\rho)\rho^{-\frac{4}{3}}=0$
@@
Due to (F2) we know $A(\rho)$ is well-defined. Easy to see $A(\rho)$ is strictly increasing. It turns out that $A(\rho)$ is also convex (see Lemma 2.6), and it is related to the internal energy $U(\rho)$ given in (7) below. Moreover, if polytropic equations of state $P(\rho)=K\rho^{\gamma}$ holds, where the parameter $\gamma>\frac{4}{3}$, then easy to check $P$ satisfies (F1) (F2) (F3), and
@@
When showing the existence of non-rotating single star, (F3) can be replaced by the assumption
@@
where $K>0$ should be large enough to prevent gravitational collapse, and thereby the Chandrasekhar mass for the model is assumed greater than $m=1$. Then we can obtain minimizers with total mass not larger than 1 via variational method. See [31, Section 2] or [5, Section 1,6,8 and Appendix].
@@
By L’Hôpital’s rule [35, 5.13 Theorem], and the integration expression of $A(s)=s\int_{0}^{s}P(t)t^{-2}dt$ in (1), we can prove that AA also satisfies (F2) and (F3) (or (F3’) if $P$ satisfies (F3’) instead of (F3), though the constant $K$ can be different).
@@
Notice $A^{\prime}(s)$ is continuous and $A^{\prime}(s)s-A(s)=P(s)$.
@@
Remark 2.5.
We claim the assumptions above can imply $A$ is convex and $A^{\prime}$ is strictly increasing. One formal way to see it, as Auchmuty and Beals suggest [5], is by taking the derivative again. Due to (F1), we know $P^{\prime}(s)$ exists almost everywhere and $P^{\prime}$ is locally integrable, (one can refer to [34, Lebesgue’s Theorem and Corollary 4 below]). Since $A^{\prime}(s)s-A(s)=P(s)$, it is conceivable $A^{\prime\prime}(s)=\frac{P^{\prime}(s)}{s}\geq 0$ a.e., which might imply $A^{\prime}(s)$ increases, and then $A(s)$ is convex [35, Chapter 5, Exercise 14]. However, it is not really sufficient. One counterexample can be $f(s)=-\Theta(s)$, where $\Theta(s)$ is Cantor-Lebesgue function, which means $f^{\prime}(s)=0$ a.e. while $f$ decreases and is not constant. Therefore, we need a more detailed analysis to prove rigorously that $A^{\prime}$ is strictly increasing.
@@
Due to (3), we can see $A^{\prime}$ is continuous and $A^{\prime}(s)>A^{\prime}(0)=0$ when $s>0$. Given $s>0$ and $h>0$, by (3) we have $A^{\prime}(s+h)-A^{\prime}(s)=\int_{s}^{s+h}P(t)t^{-2}dt+\frac{P(s+h)}{s+h}-\frac{P(s)}{s}$. Since $P$ is monotonic (strictly increasing), by second mean value theorem for definite integrals [20], we know $\exists\xi\in(0,h)$, $\int_{s}^{s+h}P(t)t^{-2}dt=P(s)\int_{s}^{s+\xi}t^{-2}dt+P(s+h)\int_{s+\xi}^{s+h}t^{-2}dt=P(s)\left(\frac{1}{s}-\frac{1}{s+\xi}\right)+P(s+h)\left(\frac{1}{s+\xi}-\frac{1}{s+h}\right)$. Then $A^{\prime}(s+h)-A^{\prime}(s)=\frac{P(s+h)-P(s)}{s+\xi}>0$, which means $A^{\prime}$ is strictly increasing on $0,\infty)$. The convexity of $A$ can refer to [35, Chapter 5, Exercise 14]
@@
Thanks to Lemma 2.6, the inverse function of $A^{\prime}$, denoted by $\left(A^{\prime}\right)^{-1}$ or $\phi$, is well defined on $0,\infty)$. Moreover, $\phi=\left(A^{\prime}\right)^{-1}$ is continuous since $A^{\prime}$ is continuous.
Another idea to show the existence of $\phi$, at least local existence, is using inverse function theorem. Assume $P(s)$ is continuously differentiable, we can know when $s>0,A^{\prime\prime}(s)$ exists and $A^{\prime\prime}(s)=\frac{P^{\prime}(s)}{s}$. However, one could run into trouble when using inverse function theorem if $P^{\prime}(s)=0$ (notice $P$ strictly increases does not mean $P^{\prime}(s)>0$, for example $f(x)=x^{3}$ at point $x=0$). But since the differentiability of $P$ is one idea to show the existence of $\nabla P(\rho)$, which is one term in the Euler-Poisson equation (EP) (or (EP’)), we can make an additional assumption of $P(\rho)$:
@@
The motivation of assuming (F4) is the following: later the minimizer $\sigma$ satisfies $\sigma=\phi\circ W_{\sigma}$ where $\sigma>0$ ($W_{\sigma}$ will be specified later), and then the pressure $P(\sigma)=P\circ\phi\circ W_{\sigma}$ in $\{\sigma>0\}$. By (F4), we know $A^{\prime\prime}(\rho)$ exists and $A^{\prime\prime}(\rho)=\frac{P^{\prime}(\rho)}{\rho}\neq 0$ is continuous if $\rho>0$, which implies $\phi=\left(A^{\prime}\right)^{-1}\in C^{1}((0,\infty))$ (one can refer to [38, Theorem 10.4.2]). It turns out this can help to show the differentiability of $\sigma$ and then of $P(\sigma)$ where $\sigma>0$, as we can see in Theorem 2.25 below.
Since actually what we want is the differentiability of $P(\sigma)$ rather than of $\sigma$ to show (EP’), another alternative assumption, which is more general but also somewhat technical, is:
@@
The advantages and applications of this more general assumption will be demonstrated later, such as in Lemma 3.1 and Theorem 2.25 (vii).
In the following sections, we assume $P(\rho)$ satisfies (F1)(F2)(F3) unless otherwise specified. We will mention (F4) or (F4’) or other assumptions when we want to use it. As a special case, if we assume the polytropic equations of state with index $\gamma>\frac{4}{3}$ holds true, one can check the assumptions (F1)(F2)(F3)(F4) or (F1)(F2)(F3)(F4’) above are satisfied automatically.
We consider a “potential admissible class” for $\rho$ and $v$ as the following:
$R\left(\mathbb{R}^{3}\right)$ $:=\left\{\left.\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\right\rvert\,\rho\geq 0,\int\rho=1\right\}$ (4)
@@
Then given $\rho$ and $v$ in such sets, McCann [31] consider energy $E(\rho,v)$ consisting of three terms:
@@
$U(\rho):=\int A(\rho(x))d^{3}x$ (7)
$G(\sigma,\rho):=\int V_{\sigma}\rho=\iint\frac{\sigma(x)\rho(y)}{|x-y|}d^{3}yd^{3}x$ (8)
$T(\rho,v):=\frac{1}{2}\int|v|^{2}\rho$ (9)
Here $A(\rho)$ is a convex function given in (1), and $U(\rho)$ is internal energy. $G$ is viewed as a symmetric bilinear form since this observation can help us to compute the variational derivative in Lemma 5.3 in subsection 5.1. Then the gravitational potential energy is defined as $G(\rho,\rho)$.
We can choose units so that the total mass of fluid is one and the gravitational constant $G=1$, then $V_{\rho}$ represents the gravitational potential of the mass density $\rho(x)$
$V_{\rho}(x):=\int\frac{\rho(y)}{|y-x|}d^{3}y$ (10)
Hence $G(\rho,\rho)$ is gravitational interaction energy.
We can also see $T(\rho,v)$ is kinetic energy.
@@
$\bar{x}(\rho):=\frac{\int x\rho(x)d^{3}x}{\int\rho(x)d^{3}x}$ (11)
is at rest. We are interested in finding minimum energy configurations subject to constraints of fixed mass ratio and angular momentum $\boldsymbol{J}$ with respect to the center of mass $\bar{x}(\rho)$. The fluid angular momentum is given by $\boldsymbol{J}(\rho,v)$:
$\boldsymbol{J}(\rho,v):=\int(x-\bar{x}(\rho))\times v\rho(x)d^{3}x$ (12)
@@
$I(\rho):=\int\rho r^{2}(x-\bar{x}(\rho))$ (13)
@@
When $\rho$ has positive mass, we have $I(\rho)>0$. If not, $I(\rho)=\int\rho(x)r^{2}(x-\bar{x}(\rho))=0$ implies $\rho(x)r^{2}(x-\bar{x}(\rho))=0$ almost everywhere in $\mathbb{R}^{3}$. Since $\mu\left(\left\{x\in\mathbb{R}^{3}\mid r^{2}(x-\bar{x}(\rho))=0\right\}\right)=0$, we have $\rho(x)=0$ almost everywhere in $\mathbb{R}^{3}$, which contradicts the fact that $\rho$ has positive mass.
A simple case is the non-rotating problem, i.e. $\boldsymbol{J}=0$, with energy $E_{0}(\rho)=U(\rho)-\frac{G(\rho,\rho)}{2}$. Properties of non-rotating minimizers are introduced in [14, 31, 5, 29].
@@
Since $\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$, by Hardy-Littlewood-Sobolev Inequality Proposition A.7 we have $V_{\rho}\in L^{12}\left(\mathbb{R}^{3}\right)$, with $\left\|V_{\rho}\right\|_{L^{12}\left(\mathbb{R}^{3}\right)}<C_{1}(\lambda)$ for some constant $C_{1}(\lambda)>0$. By Interpolation Inequality Proposition A.6 we have $\rho\in L^{\frac{12}{11}}\left(\mathbb{R}^{3}\right)$, with $\|\rho\|_{L^{\frac{12}{11}\left(\mathbb{R}^{3}\right)}}<C_{2}(\lambda)$ for some constant $C_{2}(\lambda)>0$. Finally, by Hölder’s inequality we have
$G(\rho,\rho)=\int\rho V_{\rho}\leq\|\rho\|_{L^{\frac{12}{11}\left(\mathbb{R}^{3}\right)}}\left\|V_{\rho}\right\|_{L^{12}\left(\mathbb{R}^{3}\right)}<C_{1}(\lambda)\cdot C_{2}(\lambda)\coloneq C(\lambda)$
@@
From Proposition 2.10, we see one reason why we require $\rho$ to belong not only to $L^{1}$ but also to $L^{\frac{4}{3}}$. Furthermore, the rationale behind selecting the specific exponent $\frac{4}{3}$ can also refers to McCann’s construction, which is a natural assumption by interpolation inequality Proposition A.6 if we hope $\inf\limits_{\rho\in{R}\left(\mathbb{R}^{3}\right)}E_{0}(m\rho)$ to be finite, see [5, Section 1,6 and 8], and [31, Section 2].
Consider the non-rotating minimizer $\sigma_{m}$ of $E_{0}(\rho)$ among configurations of mass $m\leq 1$, the corresponding minimum energy is finite due to the remark above. For the sake of convenience, we denote them by
@@
Intuitively we hope the minimizer should have a center of mass and finite moments of inertia, and one can see equation (EP’) will be satisfied automatically if $V_{\rho}$ has finite gradient and $x$ is the exterior point of the support of $\rho$ (denoted by spt $\rho$) as demonstrated in the proof of Theorem 2.25 (vii) in Section 3. Here support of $\rho$ is the smallest closed set carrying the full mass of $\rho$.
Therefore, we also introduce a centered subset $R_{a}\left(\mathbb{R}^{3}\right)$ of $R\left(\mathbb{R}^{3}\right)$:

### v2

@@ 1 Introduction
We consider uniformly rotating star systems, where we simply consider only the orbital revolution around each other and ignore the rotations of objects around their own axes. It is described by the reduced Euler-Poisson equation:
$\omega^{2}\tilde{\rho}(x)P_{12}(x)+\nabla P(\tilde{\rho}(x))-\tilde{\rho}(x)\nabla V_{\tilde{\rho}}(x)=0,$ (EP’)
[…]
This reduction for uniformly rotating configurations arises from the full Euler–Poisson system:
$\begin{split}\partial_{t}\rho+\nabla\cdot(\rho v)&=0,\\ \rho\partial_{t}v+\rho(v\cdot\nabla)v+\nabla P(\rho)&=\rho\nabla V,\\ \Delta V&=-4\pi\rho,\end{split}$ (EP)
which serves as a fundamental hydrodynamic model for Newtonian stellar dynamics, describing the evolution of an isolated self‑gravitating fluid. Here $\rho(x,t)\geq 0$ is the density of fluids, $t\geq 0$ is the time, $v(x,t)\in\mathbb{R}^{3}$ is the velocity, and V(x,t)-V(x,t) is the gravitational potential. The existence of single-star and multi-star systems under different settings has been established via variational methods (e.g., [4, 5, 3, 10, 11, 26, 31]) and perturbative approaches (e.g., [21, 22, 19, 36, 37, 1, 27]).
As discussed by Jang and Seok in [23] and in Subsection 2.2 below, if $\tilde{\rho}$ satisfies the reduced Euler-Poisson equation (EP’), then $\rho\left({t,y}\right)=\tilde{\rho}\left(R_{-\omega t}y\right)$, $v\left({t,y}\right)=\omega\left({-y_{2},y_{1},0}\right)^{T}$, and $V(t,y)=V_{\widetilde{\rho}}\left(R_{-\omega t}y\right)$ satisfy the full system (EP), where $R_{-\omega t}y$ is the rotation map given in Definition 2.1.
In [31], McCann used a variational method to construct uniformly rotating binary‑star solutions to (EP’). His binary stars exhibit a dual characterization: as Hamiltonian and energy minimizers under conservative constraints, and as perturbations of simpler objects like non-rotating Lane-Emden stars or relative equilibria for point masses (e.g., [23, Section 3.2]). This characterization is essential for showing that the support of the energy constrained minimizer $\rho$ is contained within a ball of a certain radius, making $\rho$ a local minimizer with respect to the topology induced by Wasserstein $L^{\infty}$ (denoted by $W^{\infty}$ hereafter) distance. As discussed in Theorem 2.22 (see also [31, Theorem 2.1]), this ensures the existence of $\nabla P(\rho)$ in $\mathbb{R}^{3}$ and validates (EP’) in the whole space. However, McCann does not thoroughly address the existence of $\nabla P(\rho)$. While McCann provides valuable insights into the $W^{\infty}$ distance, further exploration is warranted—particularly why the local minimizer in this topology has finite energy, enabling the existence of the variational derivatives. These are what we aim to refine in this paper.
@@
    Gradient Existence and Equation Derivation: We establish the existence of $\nabla P(\rho)$ for the local minimizer $\rho$, which enables a rigorous transition from the Euler-Lagrange equation to the Euler-Poisson equation. The result is formalized in Theorem 2.22 and proved in Section 3.
@@
    Energy Finiteness and Comparison of Topologies: We furthermore establish the finiteness of energy and the existence of variational derivatives for local minimizers in Lemma 5.6. We also discuss results in the alternative topology inherited from the topological vector space, where we show finite‑energy local minimizers do not exist (Proposition 5.14), yet weak local minimizers with infinite energy can be found (Proposition 5.21).
Part of the results presented in this paper are based on the author’s Master’s thesis [12]. Beyond their intrinsic mathematical interest and largely self-contained exposition, these results serve as preliminary findings and will be used in the author’s companion paper [13] on star–planet systems, which also originates from the same thesis. In [13], they serve as an alternative to direct citations of McCann’s conclusions [31], providing a more precise and rigorous foundation for the arguments therein.
@@ 2 McCann’s Construction and Results of 2-Body Systems
In this section we review McCann’s construction and results of 2-Body systems [31]. In the first subsection, we introduce some basic settings, notations, assumptions, and problem settings such as the variational formulation. In the second subsection, we focus on explaining why the general energy form $E(\rho,v)$ can be replaced by the energy with uniform rotation $E_{J}(\rho)$. In the third subsection, we present some of McCann’s nice conclusions.
@@ 2.1 Notations and problem setting
    The projection operator of $x$ onto the $x_{1}x_{2}$ plane: $P_{12}(x)=P_{12}\left(x_{1},x_{2},x_{3}\right):=\left(x_{1},x_{2},0\right)$.
@@
    $r(x):=\left(\langle x,x\rangle_{2}\right)^{\frac{1}{2}}=\sqrt{x_{1}^{2}+x_{2}^{2}}$.
@@
    Rotation map: R_θ:=(cosθ-sinθ0sinθcosθ0001)
@@
The state of a fluid may be represented by its mass density $\rho(x)\geq 0$ and velocity vector field $v(x)$ on $\mathbb{R}^{3}$. The fluid interacts with itself through Newtonian gravity; hence, we need to consider gravitational interaction energy, which will be given later. Moreover, to define internal energy, we follow Auchmuty and Beals’s assumptions [5] and first consider a general form of the pressure $P(\rho)$ as the following:
@@
    $P:[0,\infty)\rightarrow[0,\infty)$ is continuous and strictly increasing;
@@
    $\lim\limits_{\rho\rightarrow 0}P(\rho)\rho^{-\frac{4}{3}}=0$;
@@
As shown in subsection 2.2, the integral definition of $A(\rho)$ helps us obtain (EL0’). Due to (F2), we know $A(\rho)$ is well-defined. Easy to see $A(\rho)$ is strictly increasing. It turns out that $A(\rho)$ is also convex (see Lemma 2.5), and it is related to the internal energy $U(\rho)$ given in (9) below. Moreover, if the polytropic equation of state $P(\rho)=K\rho^{\gamma}$ holds, where the parameter $\gamma>\frac{4}{3}$, then easy to check $P$ satisfies (F1) (F2) (F3), and
@@
When showing the existence of non-rotating single stars, (F3) can be replaced by the assumption
@@
for some $K>0$. Usually, $K$ is supposed to be large enough to prevent gravitational collapse, and thereby the Chandrasekhar mass for the model is assumed to be greater than $m=1$. Then we can obtain minimizers with total mass not larger than 1 via variational method. See [31, Section 2] or [5, Section 1,6,8 and Appendix].
@@
By L’Hôpital’s rule [35, Theorem 5.13], and the expression of $A(s)=s\int_{0}^{s}P(t)t^{-2}dt$ in (1), we can prove that AA also satisfies (F2) and (F3) (or (F3’) if $P$ satisfies (F3’) instead of (F3), though the constant $K$ can be different).
@@
Note that $A^{\prime}(s)$ is continuous and
$A^{\prime}(s)s-A(s)=P(s)$ (4)
@@
Due to (3), we can see $A^{\prime}$ is continuous and $A^{\prime}(s)>A^{\prime}(0)=0$ when $s>0$. Given $s>0$ and $h>0$, by (3) we have
$A^{\prime}(s+h)-A^{\prime}(s)=\int_{s}^{s+h}P(t)t^{-2}dt+\frac{P(s+h)}{s+h}-\frac{P(s)}{s}$
Since $P$ is monotonic (strictly increasing), by the second mean value theorem for definite integrals [20], we know $\exists\xi\in(0,h)$,
$\int_{s}^{s+h}P(t)t^{-2}\,dt$ $=P(s)\int_{s}^{s+\xi}t^{-2}\,dt+P(s+h)\int_{s+\xi}^{s+h}t^{-2}\,dt$
$=P(s)\left(\frac{1}{s}-\frac{1}{s+\xi}\right)+P(s+h)\left(\frac{1}{s+\xi}-\frac{1}{s+h}\right)$
Then $A^{\prime}(s+h)-A^{\prime}(s)=\frac{P(s+h)-P(s)}{s+\xi}>0$, which means $A^{\prime}$ is strictly increasing on $0,\infty)$. The convexity of $A$ can refer to [35, Chapter 5, Exercise 14]. ∎
@@
Thanks to Lemma 2.5, the inverse function of $A^{\prime}$, denoted by $\left(A^{\prime}\right)^{-1}$ or $\phi$, is well-defined on $0,\infty)$. Moreover, $\phi=\left(A^{\prime}\right)^{-1}$ is continuous since $A^{\prime}$ is continuous.
Note $\nabla P(\rho)$ is one term in the Euler-Poisson equation (EP) (or (EP’)). To show the existence of $\nabla P(\rho)$, we make an additional assumption of $P(\rho)$:
@@
By (F4), we know $A^{\prime\prime}(\rho)$ exists and $A^{\prime\prime}(\rho)=\frac{P^{\prime}(\rho)}{\rho}\neq 0$ is continuous if $\rho>0$, which implies $\phi=\left(A^{\prime}\right)^{-1}\in C^{1}((0,\infty))$ (one can refer to [38, Theorem 10.4.2]). It turns out this can help to show the differentiability of $\sigma$ and then of $P(\sigma)$ where $\sigma>0$, as we can see in Theorem 2.22 below.
Since what we want is the differentiability of $P(\sigma)$ rather than of $\sigma$ to show (EP’), another alternative assumption, which is more general but also somewhat technical, is:
@@
The applications of this more general assumption will be demonstrated later, such as in Lemma 3.1 and Theorem 2.22 (vii).
In the following sections, we assume $P(\rho)$ satisfies (F1)(F2)(F3) unless otherwise specified. We will mention (F4) or (F4’) or other assumptions when we want to use them. As a special case, if we assume the polytropic equations of state with index $\gamma>\frac{4}{3}$ holds, one can check the assumptions (F1)(F2)(F3)(F4) or (F1)(F2)(F3)(F4’) above are satisfied automatically.
We consider “admissible classes” for $\rho$ and $v$ as the following:
$R\left(\mathbb{R}^{3}\right)$ $:=\left\{\left.\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\right\rvert\,\rho\geq 0,\int_{\mathbb{R}^{3}}\rho\,dx=1\right\}$ (5)
@@
We can choose units so that the total mass of the fluid is one and the gravitational constant $G=1$. Then, given $\rho$ and $v$ in the sets above, $V_{\rho}$ represents the gravitational potential of the mass density $\rho(x)$:
$V_{\rho}(x):=\int_{\mathbb{R}^{3}}\frac{\rho(y)}{|y-x|}\,dy$ (7)
Moreover, McCann [31] considers energy $E(\rho,v)$ consisting of three terms:
@@
$U(\rho):=\int_{\mathbb{R}^{3}}A(\rho(x))\,dx$ (9)
$G(\sigma,\rho):=\int_{\mathbb{R}^{3}}V_{\sigma}\rho\,dx=\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\sigma(x)\rho(y)}{|x-y|}dydx$ (10)
$T(\rho,v):=\frac{1}{2}\int_{\mathbb{R}^{3}}|v|^{2}\rho\,dx$ (11)
Here $A(\rho)$ is a convex function given in (1), and $U(\rho)$ is the internal energy. $G$ is viewed as a symmetric bilinear form since this observation can help us to compute the variational derivative in Lemma 5.3 in subsection 5.1. Then the gravitational potential energy (also called gravitational interaction energy) is defined as $G(\rho,\rho)$. We call $T(\rho,v)$ the kinetic energy.
@@
$\bar{x}(\rho):=\frac{\int_{\mathbb{R}^{3}}x\rho(x)\,dx}{\int_{\mathbb{R}^{3}}\rho(x)\,dx}$ (12)
is at rest. We are interested in finding minimum energy configurations subject to constraints of fixed (but small) mass ratio and fixed angular momentum $\boldsymbol{J}$ with respect to the center of mass $\bar{x}(\rho)$. The fluid angular momentum is given by $\boldsymbol{J}(\rho,v)$:
$\boldsymbol{J}(\rho,v):=\int_{\mathbb{R}^{3}}(x-\bar{x}(\rho))\times v\rho(x)\,dx$ (13)
@@
$I(\rho):=\int_{\mathbb{R}^{3}}\rho r^{2}(x-\bar{x}(\rho))\,dx=\int_{\mathbb{R}^{3}}\rho(x)\left((x_{1}-\bar{x}(\rho)_{1})^{2}+(x_{2}-\bar{x}(\rho)_{2})^{2}\right)\,dx$ (14)
where $r$ is given in Definition 2.1.
@@
When $\rho$ has positive mass, we have $I(\rho)>0$. If not, $I(\rho)=\int_{\mathbb{R}^{3}}\rho(x)r^{2}(x-\bar{x}(\rho))\,dx=0$ implies $\rho(x)r^{2}(x-\bar{x}(\rho))=0$ almost everywhere in $\mathbb{R}^{3}$. Since $\mu\left(\left\{x\in\mathbb{R}^{3}\mid r^{2}(x-\bar{x}(\rho))=0\right\}\right)=0$, where $\mu$ denotes the Lebesgue measure, we have $\rho(x)=0$ almost everywhere in $\mathbb{R}^{3}$, which contradicts the fact that $\rho$ has positive mass.
A simple case is the non-rotating problem, i.e., $\boldsymbol{J}=0$, with energy $E_{0}(\rho)=U(\rho)-\frac{G(\rho,\rho)}{2}$. Properties of non-rotating minimizers are introduced in [14, 31, 5, 29].
@@
Since $\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$, by Hardy-Littlewood-Sobolev Inequality [8, Theorem 1.7] we have $V_{\rho}\in L^{12}\left(\mathbb{R}^{3}\right)$, with $\left\|V_{\rho}\right\|_{L^{12}\left(\mathbb{R}^{3}\right)}<C_{1}(\lambda)$ for some constant $C_{1}(\lambda)>0$. By Interpolation Inequality [9, Section 4.2] we have $\rho\in L^{\frac{12}{11}}\left(\mathbb{R}^{3}\right)$, with $\|\rho\|_{L^{\frac{12}{11}\left(\mathbb{R}^{3}\right)}}<C_{2}(\lambda)$ for some constant $C_{2}(\lambda)>0$. Finally, by Hölder’s inequality we have
$G(\rho,\rho)=\int_{\mathbb{R}^{3}}\rho V_{\rho}\,dx\leq\|\rho\|_{L^{\frac{12}{11}\left(\mathbb{R}^{3}\right)}}\left\|V_{\rho}\right\|_{L^{12}\left(\mathbb{R}^{3}\right)}<C_{1}(\lambda)\cdot C_{2}(\lambda)\coloneq C(\lambda)$
@@
From Proposition 2.9, we see one reason why we require $\rho$ to belong not only to $L^{1}$ but also to $L^{\frac{4}{3}}$. Furthermore, the rationale behind selecting the specific exponent $\frac{4}{3}$ can also refer to McCann’s construction, which is a natural assumption by interpolation inequality [9, Section 4.2] if we hope $\inf\limits_{\rho\in{R}\left(\mathbb{R}^{3}\right)}E_{0}(m\rho)$ to be finite, see [5, Section 1, 6, and 8], and [31, Section 2].
Consider the non-rotating minimizer $\sigma_{m}$ of $E_{0}(\rho)$ among configurations of mass $m$, the corresponding minimum energy is finite due to the remark above. For the sake of convenience, we denote them by
@@
Physically speaking, if quantum mechanical effects are not considered, it is desirable for stars to be compactly supported. Therefore, we also introduce a subset $R_{a}\left(\mathbb{R}^{3}\right)$ of $R\left(\mathbb{R}^{3}\right)$: